1. Let S be the subspace of R5 defined by S = { (x1, x2, x3, x4, x5) E R5 : x1 = x2, x3 = 2x+ x}. Then the dimension of S is
  2. Let T: R-> R3 be defined as T (x,y,z) = (x+y, x-y, x+2z). Then the basis of range T is...
  3. Which of the following transformations are linear:

(i) T: R3 -> Rby T1 (u, v, w) = ( u - v + 2w, 5v - w).

(ii) T: P (R) -> R by T(P) = (integral sign from b to a) 2p(x)dx for a,b E R with a<= b

(iii) T3 : P(R) -> P(R) by T3 (P(U) = UP (U) + U



1
Expert's answer
2021-10-29T02:52:51-0400

"\begin{pmatrix}
 x_2\\
 x_2\\2x_4+x_5\\x_4\\x_5
\end{pmatrix}=x_2\begin{pmatrix}
 1\\1
 \\0\\0
\end{pmatrix}+x_4\begin{pmatrix}
 0\\0
 \\2\\1\\0
\end{pmatrix}+x_5\begin{pmatrix}
 0\\0\\1\\0\\1
\end{pmatrix}"


Matrix "S=\begin{pmatrix}
 1& 0&0\\
 1&0&0\\0&2&1\\0&1&0\\0&0&1
\end{pmatrix}"


rref of"S=\begin{pmatrix}
 1&0&0 \\
 0&1&0\\0&0&1\\0&0&0\\0&0&0
\end{pmatrix}"


The three vectors are linearly independent and they are the basis of S

Dimension of S=3


PART 2


"\begin{pmatrix}
 x+y\\x-y\\
 x+2z
\end{pmatrix}=x\begin{pmatrix}
 1\\
 1\\1\\
\end{pmatrix}+y\begin{pmatrix}
1\\-1\\0\\
\end{pmatrix}+z\begin{pmatrix}
 0 \\0
 \\2
\end{pmatrix}"



Matrix "T=\begin{pmatrix}
 1&1&0\\
 1&-1&0\\1&0&2
\end{pmatrix}"



rref of "T=\begin{pmatrix}
 1&0&0\\
 0&1&0\\0&0&1
\end{pmatrix}"


The vectors are linearly independent.

The basis of the range "T" is


"[\begin{pmatrix}
 1\\
 1\\1
\end{pmatrix}\>,\>\begin{pmatrix}
 1 \\-1
 \\0
\end{pmatrix}\>,\>\begin{pmatrix}
 0 \\0\\
 2
\end{pmatrix}]"


3, Part1


Every linear transformation is associated with a matrix


"\begin{pmatrix}
 u-v+2w\\
 5v-w
\end{pmatrix}=u\begin{pmatrix}
 1 \\
 0
\end{pmatrix}+v\begin{pmatrix}
 -1\\
 5
\end{pmatrix}+w\begin{pmatrix}
 2\\
 -1
\end{pmatrix}"



"T_1=\begin{pmatrix}
 1&-1&2\\
 0&5&-1
\end{pmatrix}"


"\implies" T1 is linear


3) Part2


"\int_a^bk2p(x)\>dx=k\int_a^b2p(x)\>dx"


"\int_a^b2p(x_1+x_2)\>dx"


"=\int_a^b2p(x_1)\>dx + \int_a^b2p(x_2)\>dx"


Therefore T2 is linear


3). Part 3



"T_3P(u_1+u_2)=(u_1+u_2)P(u_1+u_2)+(u_1+u_2)"


"T_3P(u_1)=u_1P(u_1)+u_1"


"T_3P(u_2)=u_2P(u_2)+u_2"


"\therefore \>T_3P(u_1+u_2) \mathrlap{\,/}{=}T_3P(u_1)"


"+T_3P(u_2)"


Therefore T3 is not linear